a 2 kg mass is moving with a speed of 3 m/s. What is the kinetic energy of the mass?
1
Expert's answer
2012-12-28T04:39:25-0500
T = m*v^2/2
T - the kinetic energy
m - mass
v - speed
T = 2*3^2/2 = 9 J